# AVOCADO: Adaptive Optimal Collision Avoidance driven by Opinion

### v1

@@ VII Conclusions
Simulated experiments have validated AVOCADO, showing that it is inexpensive to compute and overcomes existing approaches in terms of success rate, number of avoided collisions, computational time, trajectory length and time to reach the goal. In this sense, our solution is readily implementable in low-cost hardware devices, leaving the computational and memory resources free for any desired high-level task such as semantic mapping, autonomous tracking or package delivery. Moreover, extensive experiments with multiple robots and humans have demonstrated that AVOCADO presents zero-shot-transfer capabilities and works in scenarios with uncertainties, noise and evolving unpredictable human behaviors.
As future work, one direction is to explore the integration of AVOCADO in a higher-level planner that avoids dense crowded areas or convex obstacles. AVOCADO, since it does not have an horizon planner, sometimes enter in conflicting situations that lead to local minima or unavoidable collisions. Another research direction consists in extending AVOCADO to 3D settings with general robot dynamic models.

### v4

@@ VII Conclusions
Simulated experiments have validated AVOCADO, showing that it is inexpensive to compute and overcomes existing approaches in terms of success rate, number of avoided collisions, computational time and time to reach the goal. In this sense, our solution is readily implementable in low-cost hardware devices, leaving the computational and memory resources free for any desired high-level task such as semantic mapping, autonomous tracking or package delivery. Moreover, extensive experiments with multiple robots and humans have demonstrated that AVOCADO presents zero-shot-transfer capabilities and works in scenarios with uncertainties, noise and evolving unpredictable human behaviors.
As future work, one direction is to explore the integration of AVOCADO in a higher-level planner that avoids dense crowded areas or convex obstacles. AVOCADO, since it does not have an horizon planner, sometimes enter in conflicting situations that lead to local minima or unavoidable collisions. Another research direction consists in extending AVOCADO to general robot dynamic models, as other VO-based planners [29, 13]; and to 3-D settings. Lastly, it would be interesting to develop a method that, by just relying on onboard sensors, is able to consider how neighboring interactions are affecting the motion of the nearby agents, in order to distinguish non-cooperative behaviors stemming from the degree of cooperation of the agents from the ones enforced by their own collision avoidance goals.